# LAPO: Internalizing Reasoning Efficiency via Length-Adaptive Policy Optimization

### v1

@@ 5.2 Ablation Study on the Length Guidance for LAPO-I
Table 2: Experimental results within different length guidance for LAPO-I. Bold and underline indicate the best and second-best Pass@1 scores. w/ indicates length guidance used in LAPO-I.
@@
5.3 Ablation Study on the Statistical Metrics for Target Length Selection
@@ 5.4 Difficulty-Aware Computational Allocation
To understand the mechanisms behind LAPO’s efficiency gains, we first examine its macro-level capability to allocate computational resources in proportion to problem complexity. We investigate LAPO’s capacity for difficulty-aware reasoning length allocation by conducting systematic evaluations across mathematical reasoning benchmarks with inherent difficulty gradients: MATH Level 1-5 and AIME 2024. As shown in Figure 4, our experiments reveal a consistent positive correlation between problem complexity and reasoning length across all model variants. This data demonstrates LAPO’s fundamental capacity to proportionally scale computational investment with task demands. Specifically, both LAPO-D and LAPO-I exhibit a near-linear progression in average reasoning length from MATH Level 1 (approximately 1,000-1,200 tokens) to AIME 2024 (5,000-8,500 tokens), validating the proposed LAPO framework’s emergent capability in intrinsic difficulty awareness. The consistent scaling behaviors across different base models (DeepSeek-R1-1.5B and DeepScaleR-1.5B-Pre) and training stages underscore that LAPO develops robust, generalizable heuristics for computational resource management rather than model-specific optimizations.

### v2

@@ 6.2 Ablation Study on In-Context Guidance
Table 2: Experimental results with different length guidance for LAPO-I. Bold and underline indicate the best and second-best Pass@1 scores. w/ indicates the length guidance used in LAPO-I.
@@
6.3 Ablation on Statistical Metrics for Target Length
@@
Table 4: Robustness of LAPO-I to conflicting length instructions on MATH-500.
Method Length Constraint MATH-500
Pass@1 (%) #Tok
LAPO-I
Base N/A 86.3 2168
   +Short 500 86.0 2279
   +Long 3500 85.9 2300
LAPO-I w/ Outside
Base N/A 86.2 2251
   +Short 500 85.1 1247
   +Long 3500 86.1 2821
6.4 Analysis of Internalization
To validate that LAPO fosters genuine internalization, we stress-tested our default LAPO-I model against the w/ Outside ablation variant using adversarial Short (500 tokens) and Long (3500 tokens) length prompts. The results in Table 4 reveal a stark behavioral divergence. Our default LAPO-I remains robust, its output length staying stable around its 2200-token baseline, thus ignoring the conflicting external instructions. In contrast, the w/ Outside model is clearly influenced: its token count drops to 1247 under the Short constraint and rises to 2821 under the Long one. This comparison indicates that the placement of guidance is critical. Framing the budget as part of the model’s internal plan (inside <think>) builds a robust, internalized behavior. Framing it externally teaches superficial instruction-following. This indicates the observed robustness of LAPO-I is a direct result of our internalization mechanism.
@@ 6.5 Difficulty-Aware Computational Allocation
To understand the mechanisms behind LAPO’s efficiency gains, we examine its ability to allocate computational resources in proportion to problem complexity. We evaluate LAPO-trained models on benchmarks with clear difficulty gradients, from MATH Level 1 up to the highly complex AIME 2024. As shown in Figure 4, our models demonstrate a remarkable emergent capability for difficulty-aware resource allocation. There is a clear, near-linear positive correlation between problem complexity and the average reasoning length. On simpler problems, the models generate concise responses, while for the most challenging AIME questions, they produce extensive reasoning chains that are substantially longer than any solution observed during the training phase. This ability to extrapolate reasoning depth well beyond the bounds of their training experience is a crucial finding. It provides strong evidence that LAPO does not merely teach models to compress their outputs. Instead, it successfully imparts a generalizable principle of complexity-to-length mapping. This allows the models to dynamically and appropriately scale their computational investment when faced with novel problems of varying difficulty. The consistent scaling behavior across different base models further underscores that LAPO develops a robust, fundamental reasoning strategy rather than model-specific optimizations.
@@ A.2 Training Dynamics
A.3 Selection of Training Dataset
Table 7: Performance on the GPQA benchmark. LAPO demonstrates generalizable efficiency and accuracy gains in a non-mathematical, knowledge-intensive domain.
Method Pass@1 (%) #Tokens
Base Model: DeepSeek-R1-1.5B
Base 36.1 10297
   + LAPO-D 38.1 7596
   + LAPO-I 36.9 7235
Base Model: DeepScaleR-1.5B-Preview
Base 36.1 7667
   + LAPO-D 38.3 6176
   + LAPO-I 37.8 6154
A.4 Generalizability to Expert-Level Question Answering.
To test if LAPO’s benefits extend beyond structured mathematical reasoning, we evaluated our method on the GPQA benchmark. The results, presented in Table 7, demonstrate that LAPO’s core principles are highly generalizable.
For both base models, LAPO achieves a compelling dual improvement in accuracy and efficiency. On the DeepSeek-R1-1.5B model, LAPO-D improves Pass@1 accuracy by a significant 2.0 points while reducing token generation by 26.2%. Similarly, on the more advanced DeepScaleR-1.5B-Preview, LAPO-D boosts accuracy by 2.2 points and cuts tokens by 19.4%. The internalization stage consistently pushes efficiency further while maintaining a strong accuracy improvement over the baseline. This robust performance on a knowledge-intensive, non-mathematical task indicates that LAPO is not merely exploiting domain-specific patterns. Instead, it learns a fundamental and transferable skill: how to allocate cognitive effort efficiently for complex reasoning across different domains.